# Variational Mode Decomposition–Based Nonstationary Coherent Structure Analysis for Spatiotemporal Data

### v1

@@ 3.2.3 Parameter dependency
Figure 7: Influence of parameter $K$; time-frequency distribution of the $x$-direction velocity component at position $(x,~{}y)=(5,~{}0.5)$ for each ICS. The shade of the points represents the amplitude; $\alpha=1000$.
The time evolution of the ICS amplitudes and the time-frequency distribution of the $x$-direction velocity component at position $(x,~{}y)=(5,~{}0.5)$ for each ICS, with varying the number of ICSs ($K=4$, 5, 6, and 7), are depicted in Figs. 6 and 7, respectively. Parameter $\alpha$ was fixed at 1000. For $K=4$, the ICS corresponding to the Kármán vortex ($St_{c}=0.147$) and the ICS of $St_{c}=0.112$, representing the transient fluctuation observed for $K=5$, were not distinct; instead, they merged. For $K=6$, a new ICS corresponding to four times the frequency of the Kármán vortex ($St_{c}=0.586$) emerged. For $K=7$, in addition to the aforementioned ICSs, the ICS of $St_{c}=0.112$ split into two ICSs. As shown in Fig. 7, the variations in the instantaneous frequency for the ICSs with $St_{c}=0.289$ and $St_{c}=0.438$ were noticeable at $K=4$ and 5, but were suppressed at $K=6$, exhibiting a behavior akin to the steady frequency behavior.
Figure 8: (a) Time history of the ICS amplitudes and the (b) time-frequency distribution of the $x$-direction velocity component at position $(x,~{}y)=(5,~{}0.5)$ for each ICS. The shade of the points in (b) represents the amplitude, $K=5,~{}\alpha=1$.
Next, we investigated the dependency on parameter $\alpha$, which regulates the penalty term related to the frequency bandwidth of the ICS. Fig. 8 shows the amplitude and time-frequency distribution of the ICSs for $K=5$ and $\alpha=1$. Comparing with the results for $K=5$ and $\alpha=100$ (Fig. 6), for $\alpha=1$, i.e., a small $\alpha$ that weakens the penalty, the ICS of $St_{c}=0.112$, representing transient fluctuations, disappeared. Instead, the ICS of $St_{c}=0.003$, depicting a gradual change in the background flow, split, and the ICS of $St_{c}=0.006$ appeared. No significant changes in the obtained ICSs were observed for $\alpha=10$, 100, and 1000 (i.e., same as the results shown in Fig. 6). Thus, in this analysis, the dependency on $\alpha$ was not particularly pronounced. Note that, for $\alpha>1000$, it was not possible to obtain solutions with a sufficiently small reconstruction error (10).
@@ 4.1 Problem Setting and Input Data
In this section, the analysis of the low-Reynolds-number flow around a pitching airfoil is reported. The analyzed data were obtained from a publicly accessible database Towne2023 . These data comprised a two-dimensional flow field around a flat-plate airfoil with periodic pitching motion. The freestream Reynolds number based on the airfoil chord length was 100. The angle of attack of the airfoil oscillates according to the equation:
@@
According to Ref. Towne2023 , vortex shedding occurs at a frequency of $St=0.2418$ even without pitching at an angle of attack of 30 deg. This flow field exhibits primarily two fluctuating phenomena: vortex shedding and fluctuation due to pitching, resulting in a quasi-periodic flow field. This analysis aims to investigate the mechanism through which these phenomena are captured by VMD-NCS analysis.
The mean-subtracted vorticity field around a flat-plate airfoil was employed as the input datasets for the VMD-NCS analysis. Each snapshot had equally spaced $599\times 299$ grid points, and the number of snapshots was 1001. The time interval between the snapshots was $\Delta t=0.1$. For further details of the datasets, please refer to Ref. Towne2023 .
@@ 4.2.1 Intrinsic coherent structure frequency distribution
The center frequencies and time-averaged amplitudes of the ICS with varying parameters $K$ and $\alpha$ are shown in Fig. 9. The amplitudes are normalized so that the largest amplitude is unity in the figure. For $K=2$ (Fig. 9(a)), a similar distribution is observed for all $\alpha$ cases. The center frequencies of the two ICSs were $St_{c}=0.063$ and $St_{c}=0.258$. These frequencies were close to the pitching frequency ($f_{p}=0.05$) and the vortex shedding frequency ($St=0.2418$), respectively, indicating the successful extraction of phenomena related to their respective variations. As the number of the ICSs increased, the dependence on $\alpha$ appeared. For $K=4$, there was little difference between the results for $\alpha=20$ and 200, where, besides the ICS corresponding to the pitching and vortex shedding frequencies, the second and third harmonics of the vortex shedding frequency ($St_{c}=0.461$ and $St_{c}=0.714$, respectively) were extracted. For the smallest value (corresponding to the weakest penalty), $\alpha=2$, a different ICS emerged around the vortex shedding frequency, resulting in a different frequency distribution. Upon further increasing the $K$, the ICS associated with the main vortex shedding frequency and its second harmonic were further decomposed, generating ICSs at frequencies plus and minus the pitching frequency, as shown in Fig. 9(d). At $\alpha=2000$, the reconstruction error (10) increased for all $K$ values, and a convergent solution could not be obtained.
@@
Fig. 10 illustrates the frequencies and mean amplitudes of the DMD modes. Here, the TLS-DMD algorithm Hemati2017 was employed for the DMD, and the mean amplitudes of the DMD modes were computed following Kou et al. Kou2017 . For comparison, the ICSs for $K=14$ and $\alpha=200$ are also included in the figure. From this figure, it is evident that the distribution of the center frequencies and mean amplitudes of ICSs for $K=14$ closely resembled those of the DMD modes. Specifically, the modes of the vortex shedding frequencies plus and minus the pitching frequencies appeared, and their amplitudes closely matched between the ICSs and DMD modes. This tendency of the ICS obtained at a large $K$ to represent more single-frequency variations (i.e., similar to the case with the DMD modes) was also observed in the analysis described in Section 3.
In summary, the center frequencies of the ICS obtained from this flow field exhibited similarities to those obtained using the DMD method for large $K$ values, whereas for small $K$ values, the phenomenon of vortex shedding influenced by wing pitching motion was captured as a single ICS.
@@
Figure 11: The temporal evolution of the expansion coefficients of the (a) first, (b) third, (c) fifth, (d) seventh, and (e) ninth POD modes (left); the MVMD modes obtained by decomposing the expansion coefficients (middle); and the frequency distribution of the MVMD modes (right).
Next, the results of the VMD-NCS analysis obtained with $K=4$ and $\alpha=200$ are discussed. The temporal evolutions of the POD expansion coefficients, $\tilde{q}_{i}(t)$, and the MVMD modes, $u_{k,i}(t)$, as well as their frequency distributions, are shown in Fig. 11. From this figure, similar to the results in Section 3, it is evident that each POD coefficient exhibits multiple frequency-band fluctuations. For instance, the temporal evolution of the coefficients of the fifth POD mode exhibited a combination of variations corresponding to both the pitching and vortex shedding frequencies (Fig. 11(c)). Moreover, the variations corresponding to the pitching frequency were also contained in the first and fifth POD modes, and those corresponding to the vortex shedding frequency were present in the third, fifth, and seventh POD modes. Therefore, to extract the fluid phenomena associated with each frequency band, it is necessary to isolate and superimpose the frequency components dispersed among multiple POD modes. The VMD-NCS analysis achieves this by decomposing the POD coefficients using MVMD. From the frequency distributions, it can be confirmed that the temporal variations of the MVMD modes, which constitute the ICSs, primarily exhibited frequency components around their corresponding center frequencies ($St_{c}=0.56,~{}0.236,~{}0.461,~{}0.714$).
@@ 4.2.3 Temporal behavior of the intrinsic coherent structure ($K=4$, $\alpha=200$)
Figure 12: Time history of the ICS amplitudes.
Figure 13: Spatial distribution of the vorticity of the instantaneous flow field (left), the ICS of $St_{c}=0.056$ (middle), and the ICS of $St_{c}=0.236$ (right); (a), (b), (c), and (d) correspond to the times indicated in Fig. 12; $\theta$ is the phase in Equation (21), i.e., $\theta=2\pi f_{p}t$.
Figure 14: Spatial distribution of the DMD modes of $St=0.05$ (left) and $St=0.238$ (right); (a), (b), (c), and (d) correspond to the time indicated in Fig. 12; $\theta$ is the phase in Equation (21), i.e., $\theta=2\pi f_{p}t$.
Fig. 12 shows the time history of the ICS amplitudes, and Fig. 13 illustrates the spatial distribution of the instantaneous flow fields and ICSs at each phase (a)–(d) depicted in Fig. 12. By observing the time variation of the amplitude of the ICS corresponding to vortex shedding in Fig. 12 ($St_{c}=0.236$), we note that the amplitude reached its maximum at the minimum angle of attack (25 deg) and achieves its minimum at the maximum angle of attack (35 deg). This indicates that the fluctuations attributed to vortex shedding were significantly influenced by the airfoil’s pitching motion, weakened with pitching-up and strengthened with pitching-down. Conversely, the amplitude of the ICS ($St_{c}=0.056$) corresponding to the frequency of the pitching motion was not significantly affected by the vortex shedding.
From the spatial distribution shown in the right column of Fig. 13, it can be observed that the ICS of $St_{c}=0.236$ effectively described the influence of the pitching motion on the spatial patterns of the vortex shedding phenomena. For instance, comparing Figs. 13(a) and (c), at the same angle of attack of 30 deg, it is clear that the spatial pattern of the ICS ($St_{c}=0.236$) differs, indicating greater fluctuation near the airfoil during pitching-down (a) compared with that during pitching-up (c).
To compare the ICSs with the DMD modes, Fig. 14 illustrates the spatial distribution of the DMD modes. Here, the spatial distribution of each DMD mode at time $t=t_{j}$, $\psi_{DMD}(\bm{x},t_{j})$, is defined as the term of the respective mode when the flow field, $q(\bm{x},t_{j})$, is represented as a superposition of the DMD modes. It was calculated using the following equation:
$\psi_{DMD}(\bm{x},t_{j})=2{\rm{Re}}\left(c(t_{j})\phi_{DMD}(\bm{x})\right)$ (22)
Here, $c(t_{j})$ represents the expansion coefficient at time $t=t_{j}$, $\phi_{DMD}$ represents the DMD eigenvector, and ${\rm Re}(\cdot)$ denotes the real part of a complex number. As the spatial distribution of the DMD eigenvector, $\phi_{DMD}$, remains unchanged over time, it is evident from the right column of Fig. 14 that the spatial distribution of the DMD modes corresponding to the vortex shedding frequency was not influenced by the phase of the pitching motion. As demonstrated in Fig. 13, an advantage of the ICS obtained by the VMD-NCS analysis is its capability to represent spatial-distribution changes in time, which cannot be expressed by a single DMD mode.
@@ 5 Conclusions
Validation targeting the transient growth phenomenon around a square cylinder revealed that the VMD-NCS analysis method can extract spatiotemporal patterns (as ICSs) representing not only the periodic shedding of the Kármán vortices but also the transitional phenomena from steady flow to periodic flow. Additionally, the extraction of ICS depicting the gradual temporal change of the background flow was demonstrated. A close examination of the temporal variation in the amplitudes of these ICSs revealed observations, for example, an increase in the amplitude of the transient phenomena resulted in vortex generation along with a decrease in the amplitude of the background flow. Moreover, during the transient growth phase of vortex shedding, the ICS represented the temporal variation in the spatial distribution, initially showing fluctuations in the far-wake region and gradually intensifying fluctuations in the near-wake region. In the analysis of the flow around the pitching airfoil, we examined the mechanism through which the VMD-NCS analysis method represents quasi-periodic flow fields resulting from the airfoil’s periodic pitching motion and the vortex shedding in the wake region. The results indicated that the size and amplitude of vortices in the wake region were influenced by the pitching motion, i.e., weakened and strengthened during pitching-up and pitching-down, respectively, represented as temporal changes in a single ICS. In all these analyses, the ICS derived through VMD-NCS demonstrated a ‘flexible’ attribute Noack2016 , allowing the capture of temporal changes in spatial patterns as a single ICS, a capability not typically achievable through conventional modal analysis techniques, such as POD and DMD.
@@
Thus, the VMD-NCS analysis serves as an effective method for extracting spatiotemporal patterns of nonstationary phenomena, which is difficult to achieve using conventional methods. The ICS extracted using the VMD-NCS analysis method is ‘flexible,’ evolving with time in terms of spatial distribution and maintaining temporal coherence, thus improving the comprehension of phenomena. However, the dependency of results on the parameters, i.e., $K$ and $\alpha$, poses a limitation. Future studies will delve further into understanding the parameter dependence and exploring methods for determining appropriate parameters.

### v2

@@ 3.2.3 Parameter dependency
Figure 8: Influence of parameter $K$; time-frequency distribution of the $x$-direction velocity component at position $(x,~{}y)=(5,~{}0.5)$ for each ICS. The shade of the points represents the amplitude. $\alpha=1000$.
The time evolution of the ICS amplitudes and the time-frequency distribution of the $x$-direction velocity component at position $(x,~{}y)=(5,~{}0.5)$ for each ICS, with varying the number of ICSs ($K=4$, 5, 6, and 7), are depicted in Figs. 7 and 8, respectively. Parameter $\alpha$ was fixed at 1000. For $K=4$, the ICS corresponding to the Kármán vortex ($St_{c}=0.147$) and the ICS of $St_{c}=0.112$, representing the transient fluctuation observed for $K=5$, were not distinct; instead, they merged. For $K=6$, a new ICS corresponding to four times the frequency of the Kármán vortex ($St_{c}=0.586$) emerged. For $K=7$, in addition to the aforementioned ICSs, the ICS of $St_{c}=0.112$ split into two ICSs. As shown in Fig. 8, the variations in the instantaneous frequency for the ICSs with $St_{c}=0.289$ and $St_{c}=0.438$ were noticeable at $K=4$ and 5, but were suppressed at $K=6$, exhibiting a behavior akin to the steady frequency behavior.
Figure 9: (a) Time history of the ICS amplitudes and the (b) time-frequency distribution of the $x$-direction velocity component at position $(x,~{}y)=(5,~{}0.5)$ for each ICS. The shade of the points in (b) represents the amplitude. $K=5,~{}\alpha=1$.
Next, we investigated the dependency on parameter $\alpha$, which regulates the penalty term related to the frequency bandwidth of the ICS. Fig. 9 shows the amplitude and time-frequency distribution of the ICSs for $K=5$ and $\alpha=1$. When compared to the results for $K=5$ and $\alpha=100$ (Fig. 7), setting $\alpha=1$ which represents a weaker penalty, leads to the disappearance of the ICS with $St_{c}=0.112$ which captured the transient fluctuations. Instead, a new ICS with $St_{c}=0.006$ emerged. No significant changes in the obtained ICSs were observed for $\alpha=10$, 100, and 1000 (i.e., same as the results shown in Fig. 7). Thus, in this analysis, the dependency on $\alpha$ was not particularly pronounced. Note that, for $\alpha>1000$, it was not possible to obtain solutions with a sufficiently small reconstruction error (10).
@@ 4.1 Problem Setting and Input Data
In this section, the analysis of the low-Reynolds-number flow around a pitching airfoil is reported. The analyzed data were obtained from a publicly accessible database [42]. The data comprised a two-dimensional flow field around a flat-plate airfoil with periodic pitching motion. The freestream Reynolds number based on the airfoil chord length was 100. The angle of attack of the airfoil oscillates according to the equation:
@@
According to Ref. [42], vortex shedding occurs at a frequency of $St=0.2418$ even without pitching at an angle of attack of 30 deg. This flow field exhibits primarily two fluctuating phenomena: vortex shedding and fluctuation due to pitching, resulting in a quasi-periodic flow field. This analysis aims to investigate the mechanism through which these phenomena are captured by VMD-NCS analysis.
The mean-subtracted vorticity field around a flat-plate airfoil was employed as the input datasets for the VMD-NCS analysis. Each snapshot had equally spaced $599\times 299$ grid points, and the number of snapshots was 1001. The time interval between the snapshots was $\Delta t=0.1$. For further details of the datasets, please refer to Ref. [42].
@@ 4.2.1 Intrinsic coherent structure frequency distribution
The center frequencies and time-averaged amplitudes of the ICS with varying parameters $K$ and $\alpha$ are shown in Fig. 10. The amplitudes are normalized so that the largest amplitude is unity in the figure. For $K=2$ (Fig. 10(a)), a similar distribution is observed for all $\alpha$ cases. The center frequencies of the two ICSs were $St_{c}=0.063$ and $St_{c}=0.258$. These frequencies were close to the pitching frequency ($f_{p}=0.05$) and the vortex shedding frequency ($St=0.2418$), respectively, indicating the successful extraction of phenomena related to their respective variations. As the number of the ICSs increased, the dependence on $\alpha$ appeared. For $K=4$, there was little difference between the results for $\alpha=20$ and 200, where, besides the ICS corresponding to the pitching and vortex shedding frequencies, the second and third harmonics of the vortex shedding frequency ($St_{c}=0.461$ and $St_{c}=0.714$, respectively) were extracted. For the smallest value (corresponding to the weakest penalty), $\alpha=2$, a different ICS emerged around the vortex shedding frequency, resulting in a different frequency distribution. Upon further increasing the $K$, the ICS associated with the main vortex shedding frequency and its second harmonic were further divided, generating ICSs at frequencies plus and minus the pitching frequency, as shown in Fig. 10(d). At $\alpha=2000$, the reconstruction error (10) increased for all $K$ values, and a convergent solution could not be obtained.
@@
Fig. 11 illustrates the frequencies and mean amplitudes of the DMD modes. Here, the TLS-DMD algorithm [21] was employed for the DMD, and the mean amplitudes of the DMD modes were computed following Kou et al. [43]. For comparison, the ICSs for $K=14$ and $\alpha=200$ are also included in the figure. From this figure, it is evident that the distribution of the center frequencies and mean amplitudes of ICSs for $K=14$ closely resembled those of the DMD modes. Specifically, the modes of the vortex shedding frequencies plus and minus the pitching frequencies appeared, and their amplitudes closely matched between the ICSs and DMD modes. This tendency of the ICS obtained at a large $K$ to represent more single-frequency variations (i.e., similar to the case with the DMD modes) was also observed in the analysis described in Section 3.
In summary, the center frequencies of the ICS obtained from this flow field exhibited similarities to those obtained using the DMD method for large $K$ values. However, the key advantage of ICS became evident for small $K$ values, and the phenomenon of vortex shedding influenced by wing pitching motion was captured as a single ICS as shown in Sec. 4.2.3.
@@
Figure 12: The temporal evolution of the expansion coefficients (left) and MVMD modes obtained by decomposing the expansion coefficients (right) for the (a) first, (b) third, (c) fifth, (d) seventh, and (e) ninth POD modes.
Figure 13: The frequency distribution of the expansion coefficients of the (a) first, (b) third, (c) fifth, (d) seventh, and (e) ninth POD modes and their MVMD modes.
Next, the results of the VMD-NCS analysis obtained with $K=4$ and $\alpha=200$ are discussed. Fig. 12 shows the temporal evolution of the POD expansion coefficients $\tilde{q}_{i}(t)$ and the MVMD modes $u_{k,i}(t)$. Additionally, Fig. 13 presents their frequency distributions. From this figure, similar to the results in Section 3, it is evident that each POD coefficient exhibits multiple frequency-band fluctuations. For example, the temporal evolution of the fifth POD mode coefficients showed a combination of variations related to both pitching and vortex shedding frequencies, as seen in Figs. 12(c) and 13(c). Additionally, variations associated with the pitching frequency appeared not only in the fifth but also in the first POD mode. Similarly, variations linked to the vortex shedding frequency were present across multiple modes: the third, fifth, and seventh POD modes. Therefore, to extract the fluid phenomena associated with each frequency band, it is necessary to isolate and superimpose the frequency components dispersed among multiple POD modes. The VMD-NCS analysis achieves this by decomposing the POD coefficients using MVMD. From the frequency distributions (Fig. 13), it can be confirmed that the temporal variations of the MVMD modes, which constitute the ICSs, primarily exhibited frequency components around their corresponding center frequencies ($St_{c}=0.56,~{}0.236,~{}0.461,~{}0.714$).
@@ 4.2.3 Temporal behavior of the intrinsic coherent structure ($K=4$, $\alpha=200$)
Figure 14: Time history of the ICS amplitudes. (a), (b), (c), and (d) in the figure indicate the times when $a(t)=30^{\circ},~{}25^{\circ},~{}30^{\circ}$, and $35^{\circ}$, respectively.
Figure 15: Spatial distribution of the vorticity of the instantaneous flow field, the ICS of $St_{c}=0.056$, the ICS of $St_{c}=0.236$, and the DMD mode of $St=0.238$; (a), (b), (c), and (d) correspond to the times indicated in Fig. 14; $\theta$ is the phase in Equation (21), i.e., $\theta=2\pi f_{p}t$.
Fig. 14 shows the time history of the ICS amplitudes, and Fig. 15 illustrates the spatial distribution of the instantaneous flow fields and ICSs at each phase (a)–(d) depicted in Fig. 14. By observing the time variation of the amplitude of the ICS corresponding to vortex shedding in Fig. 14 ($St_{c}=0.236$), we note that the amplitude reached its maximum at the minimum angle of attack (25 deg) and achieves its minimum at the maximum angle of attack (35 deg). This indicates that the fluctuations attributed to vortex shedding were influenced by the airfoil’s pitching motion, weakened with pitching-up and strengthened with pitching-down. Conversely, the amplitude of the ICS ($St_{c}=0.056$) corresponding to the frequency of the pitching motion was not significantly affected by the vortex shedding.
From the spatial distribution shown in Fig. 15, it can be observed that the ICS of $St_{c}=0.236$ described the influence of the pitching motion on the spatial patterns of the vortex shedding phenomena. For instance, comparing Figs. 15(a) and (c), at the same angle of attack of 30 deg, it is clear that the spatial pattern of the ICS ($St_{c}=0.236$) differs, indicating greater fluctuation near the airfoil during pitching-down (a) compared with that during pitching-up (c).
For comparison with the DMD mode, the spatial distributions of the DMD mode ($St=0.238$) are shown in the fourth column of Fig. 15. Here, the spatial distribution of each DMD mode at time $t=t_{j}$, denoted as $\psi_{DMD}(\bm{x},t_{j})$, is defined as its contribution to the flow field $q(\bm{x},t_{j})$ when $q(\bm{x},t_{j})$ is expressed as a superposition of all DMD modes. It was calculated using the following equation:
$\psi_{DMD}(\bm{x},t_{j})=2{\rm{Re}}\left(w(t_{j})\phi_{DMD}(\bm{x})\right)$ (22)
Here, $w(t_{j})$ represents the expansion coefficient at time $t=t_{j}$, $\phi_{DMD}$ represents the DMD eigenvector, and ${\rm Re}(\cdot)$ denotes the real part of a complex number. As the spatial distribution of the DMD eigenvector, $\phi_{DMD}$, remains unchanged over time, it is evident that the spatial distribution of the DMD modes corresponding to the vortex shedding frequency ($St=0.238$) was not influenced by the phase of the pitching motion. As demonstrated in Fig. 15, an advantage of the ICS obtained by the VMD-NCS analysis is its capability to represent spatial-distribution changes in time, which cannot be expressed by a single DMD mode.
@@ 5 Conclusions
Validation targeting the transient growth phenomenon around a square cylinder revealed that the VMD-NCS analysis method can extract spatiotemporal patterns (as ICSs) representing not only the periodic shedding of the Kármán vortices but also the transitional phenomena from steady flow to periodic flow. Additionally, the extraction of ICS depicting the gradual temporal change of the background flow was demonstrated. A close examination of the temporal variation in the amplitudes of these ICSs revealed observations, for example, an increase in the amplitude of the transient phenomena resulted in vortex generation along with a decrease in the amplitude of the background flow. Moreover, during the transient growth phase of vortex shedding, the ICS represented the temporal variation in the spatial distribution, initially showing fluctuations in the far-wake region and gradually intensifying fluctuations in the near-wake region. In the analysis of the flow around the pitching airfoil, we examined the mechanism through which the VMD-NCS analysis method represents quasi-periodic flow fields resulting from the airfoil’s periodic pitching motion and the vortex shedding in the wake region. The results indicated that the size and amplitude of vortices in the wake region were influenced by the pitching motion, i.e., weakened and strengthened during pitching-up and pitching-down, respectively, represented as temporal changes in a single ICS. In all these analyses, the ICS derived through VMD-NCS demonstrated a ‘flexible’ attribute [32], allowing the capture of temporal changes in spatial patterns as a single ICS, a capability not typically achievable through conventional modal analysis techniques, such as POD and DMD.
@@
Thus, the VMD-NCS analysis serves as an effective method for extracting spatiotemporal patterns of nonstationary phenomena, which is difficult to achieve using conventional methods. The ICS extracted using the VMD-NCS analysis method is ‘flexible,’ evolving with time in terms of spatial distribution and maintaining temporal coherence, thus improving the comprehension of phenomena. However, the dependency of results on the parameters, i.e., $K$ and $\alpha$, remains a limitation. Future studies will delve further into understanding the parameter dependence and exploring methods for determining appropriate parameters.